OpenStudy (anonymous):

i need help pleace write an equation of the line with the given slope and y-intercept slope 7/6,y-intercept:3 a.y=7/6x+3 b.y=-7/6x+3 c.y=7/6x-3 d.y=6/7x+3

OpenStudy (texaschic101):

point slope formula is y = mx + b. m = slope and b = y intercept. Now sub the information you were given into your formula.

OpenStudy (anonymous):

im not good at any of this sorry

OpenStudy (texaschic101):

it ok...that is why you are here on this site.

OpenStudy (anonymous):

true i just look for a or b cause im taking online

OpenStudy (texaschic101):

ok.....m (which is the slope) = 7/6....and b (which is the y intercept) = 3. The formula is y = mx + b , now sub in and the answer is y = 7/6x + 3

OpenStudy (anonymous):

sweet thanks

OpenStudy (texaschic101):

your quite welcome....have a nice day :)

OpenStudy (anonymous):

u to